# Long distance propagation of wave beams in paraxial regime

### v1

@@ Abstract
This paper concerns the propagation of high frequency wave-beams in highly turbulent atmospheres. Using a paraxial model of wave propagation, we show in the long-distance weak-coupling regime that the wavefields are approximately described by a complex Gaussian field whose scintillation index is unity. This provides a model of the speckle formation observed in many practical settings. The main step of the derivation consists in showing that closed-form moment equations in the Itô-Schrödinger regime are still approximately satisfied in the paraxial regime. The rest of the proof is then an extension of results derived in [5].
@@ 1 Introduction
$\Big{(}\partial^{2}_{z}+\Delta_{x}+k_{0}^{2}n^{2}(z,x)\Big{)}p(z,x)=-u_{0}(x)\delta_{0}(z)\,,$ (1)
@@
with $0<\theta\ll 1$. As in [5], we consider the weak coupling regime $0<\varepsilon\ll 1$ in two sub-regimes, the kinetic regime when $\eta=1$ and the diffusive regime when $\eta\ll 1$. That the latter regime corresponds to long-distance propagation (optically thick media) may be obtained from the fact that distance of propagation times random strength (proportional to $\nu^{2}$) is $\eta/\varepsilon\times\varepsilon/\eta^{3}=\eta^{-2}\gg 1$. We will see below that $\varepsilon$ models the width of the incident beam $u_{0}$. For technical reasons as in [5], we need a large scale separation between $\varepsilon$ and $\eta$ in the diffusive regime and assume that $\eta=(\log|\log\varepsilon|)^{-1}$.
@@
The above equation is deduced from (4) by formally replacing $\theta^{-\frac{1}{2}}\nu(\frac{z}{\theta},x)\mathrm{d}z$ by its central limit approximation $\mathrm{d}B(z,x)$, where $B(z,x)$ is a mean-zero Gaussian process satisfying $\mathbb{E}[B(z,x)B(z^{\prime},y)]=\min(z,z^{\prime})R(x-y)$ with spatial covariance $R(x)$ defined in (10) below. The product $u^{\varepsilon}\mathrm{d}B$ should be interpreted in a standard Itô form [10, 11].
That the complex circular Gaussian conjecture holds for $u^{\varepsilon}$ in the limit $\varepsilon\to 0$ was shown recently in [5]. Careful moment estimates of $u^{\theta}$ allow us to extend such a result to the paraxial model of wave propagation.
@@ 2 Description of the wavefield, random medium, and moments
For simplicity, we set $12k_{0}=$ 1 since this (rescaled) parameter does not have any qualitative influence, and assume that the amplitude $\nu$ is appropriately rescaled so that the paraxial equation becomes
@@
$\partial_{z}\hat{u}^{\theta}=-\frac{i\eta}{\varepsilon}|\xi|^{2}\hat{u}^{\theta}+\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\int\limits_{\mathbb{R}^{d}}\hat{\nu}\Big{(}\frac{\eta z}{\varepsilon\theta},k\Big{)}\hat{u}^{\theta}(z,\xi-k)\frac{\mathrm{d}k}{(2\pi)^{d}},\quad\hat{u}^{\theta}(0,\xi)=\hat{u}^{\theta}_{0}(\xi)\,.$
For a collection of points $(X,Y)=(x_{1},\cdots,x_{p},y_{1},\cdots,y_{q})\in\mathbb{R}^{(p+q)d}$, define the $p+q$th moment of the wavefield $u^{\theta}$ at fixed $z$ by
@@
Gaussian random potential.
@@
We assume that ^$\hat{\mathfrak{C}}(z,\cdot)$ is integrable in $z$, smooth in $k$ and rapidly decaying for large values of $z$. For concreteness and to simplify some derivations, we assume that for some $\tau_{R}>0$ and $\langle k\rangle:=\sqrt{1+|k|^{2}}$, we have
$\int_{\mathbb{R}^{d}}\langle{k}\rangle^{{d_{0}+3}}\hat{\mathfrak{C}}(z,k)\mathrm{d}k\leq\lambda_{1}(d_{0})e^{-\tau_{R}|z|},$ (9)
where we assume that $d_{0}\geq d$. Except for improved regularity results obtained when $d_{0}\geq d$ in Lemma 4.7 (and Lemma 4.2) and for better stochastic continuity properties in Theorem 3.3, there is no loss in generality to assume that $d_{0}=d$, the lateral spatial dimension. We assume the standard symmetries for correlation functions: $\mathfrak{C}(z,x)=\mathfrak{C}(-z,x)=\mathfrak{C}(z,-x)$ and denote by $R(x)$ the lateral spatial covariance
@@
We next assume that $R\in{\mathsf{L}}^{1}(\mathbb{R}^{d})\cap{\mathsf{L}}^{\infty}(\mathbb{R}^{d})$ with the Fourier transform $\hat{R}(k)=\hat{R}(-k)\leq\hat{{\mathsf{R}}}(k)=\hat{{\mathsf{R}}}(|k|)$ for some radially symmetric function $\hat{{\mathsf{R}}}$.
@@
with the Fourier transform ^$\hat{\mathfrak{C}}^{\theta}(s,k)$ defined in a similar manner.
Assumptions on the source.
@@
uniformly in $\varepsilon\in(0,1]$.
[…]
Spatially rescaled random vector.
@@ 3 Main results
The limiting approximations for $\psi^{\theta}$ as $\theta\to 0$ in the kinetic and diffusive regimes are the same as for the Itô-Schrödinger model (as $\varepsilon\to 0$) treated in [5]. The limits are described by second-order moments, which we now recall.
@@
$M_{1,1}(z,r,x,y)=\begin{cases}|u_{0}(r)|^{2}\mathcal{Q}(y-x,0),\quad&\beta>1\\ \displaystyle\int\limits_{\mathbb{R}^{d}}e^{i\xi\cdot(r-r^{\prime})}|u_{0}(r^{\prime})|^{2}\mathcal{Q}(y-x,\xi)\frac{\mathrm{d}\xi\mathrm{d}r^{\prime}}{(2\pi)^{d}},\quad&\beta=1\,,\end{cases}$ (16)
@@
$M_{1,1}(z,r,x,y)=\begin{cases}|u_{0}(r)|^{2}e^{\frac{z}{2}(y-x)^{\mathrm{t}}\Xi(y-x)},\quad&\beta>1\\[5.69054pt] e^{\frac{1}{8}z(y-x)^{\mathrm{t}}\Xi(y-x)}e^{-\frac{3i}{4z}(y-x)\cdot r}G(z^{3},r)\ast[e^{\frac{3i}{4z}(y-x)\cdot r}|u_{0}(r)|^{2}],\quad&\beta=1\,,\end{cases}$ (18)
@@
In the kinetic regime $\eta=1$, $\Phi^{\theta}(z,r,X)\Rightarrow\Phi(z,r,X)$ in distribution as $\theta\to 0$ where $\Phi(z,r,X)$ is a random vector with elements $\{\phi_{j}\}_{j=1}^{N}$ satisfying $\mathbb{E}\phi_{j}=M_{1,0}(z,r,x_{j})$ and $Z=\Phi-\mathbb{E}\Phi$ a mean zero circular complex Gaussian random vector with elements $\{Z_{j}\}_{j=1}^{N}$ satisfying
$\mathbb{E}[Z_{j}Z^{\ast}_{l}]=M_{1,1}(z,r,x_{j},x_{l})-M_{1,0}(z,r,x_{j})M_{0,1}(z,r,x_{l}),\quad\mathbb{E}[Z_{j}Z_{l}]=0\,.$ (20)
@@
$\sup\limits_{0\leq s\leq z}\mathbb{E}|\phi^{\theta}(s,r,x+h)-\phi^{\theta}(s,r,x)|^{2n}\leq C(n,\alpha,z)|h|^{2n\alpha}$ (22)
for $\|h\|_{2}<1$ uniformly in $\theta\in(0,\frac{1}{2}]$. Here, $\alpha\in(0,1]$ in the kinetic regime and $\alpha\in(0,1)$ in the diffusive regime. Let $n$ be such that $2n\alpha\geq d+2n\alpha_{-}$, $\alpha_{-}\in(0,\alpha)$. Then the process $\phi^{\theta}(z,r)$ is tight on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ and there exists a Hölder continuous version of $\phi^{\theta}(z,r)$ on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$.
@@
For $z$ and $r$ fixed and under the hypotheses of Theorem 3.3, the process $\phi^{\theta}(z,r,\cdot)$ converges in distribution as probability measures on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ to its unique limit, which is characterized by limits of finite dimensional distributions shown to converge in Theorems 3.1 and 3.2 in the kinetic and diffusive regimes.
@@
Scaling in applications.
We briefly revisit the scaling (2) and its practical relevance. There, $x$ was measured in terms of the smallest scale of turbulence, let us call it $l_{0}$, typically taken in practice to be in the $mm$ range. This means that medium fluctuations should be understood as $\nu=\nu\big{(}\frac{z}{l_{0}},\frac{x}{l_{0}}\big{)}$. With $Z$ a dimension-less parameter characterizing longitudinal distance of propagation, $w_{0}$ a parameter describing the width of the source, and $\sigma$ a covariance parameter for the random fluctuations $\nu=\sigma\tilde{\nu}$, we have the scaling and paraxial model after neglecting backscattering:
@@
A typical wavelength used in practice is $\lambda\approx 10^{-6}m$ and corresponds to $k_{0}=10^{6}m^{-1}$. Under weak turbulence assumptions [9, 11], the variations in the refractive index are consistent with $\sigma=10^{-4}$ with a fluctuation lengthscale $l_{0}=10^{-3}m$. Thus, $\theta=10^{-3}$ and $\sigma^{2}l_{0}^{2}k_{0}^{2}=10^{-2}$.
We thus observe that we reach the kinetic regime $\eta=1$ for a propagation distance of order $Zm=10^{2}m$. We are in the scintillation regime when $\varepsilon\ll 1$ or $l_{0}\ll w_{0}$, which is the case for the beam widths of a (quite a) few millimeters to meters found in applications [25, 28, 27, 1].
The diffusive regime is reached for $\eta\ll 1$, corresponding to $Z\gg 10^{2}$, with values of $Zm$ in the range $10^{3}-10^{4}m$ typical in applications. For $Z=10^{4}$ so that $\eta=10^{-1}$ for concreteness, the proposed theory has a domain of validity for beam width $w_{0}=10^{-3}/\varepsilon\gg 10^{-3}/\eta=10^{-2}m$, i.e., for beams with a diameter at least in the tens of centimeters to meter scale.
@@ 4.1 Phase compensated wave field and preparatory calculations
$\partial_{z}\tilde{u}^{\theta}=\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\int\limits_{\mathbb{R}^{d}}\hat{\nu}\Big{(}\frac{\eta z}{\varepsilon\theta},k\Big{)}\tilde{u}^{\theta}(z,\xi-k)e^{\frac{i\eta z}{\varepsilon}g(\xi,k)}\frac{\mathrm{d}k}{(2\pi)^{d}},\quad\tilde{u}^{\theta}(0,\xi)=\hat{u}^{\theta}_{0}(\xi)\,,$ (25)
@@
$L^{\theta}(s,\nu_{z})\psi=\int\limits_{\mathbb{R}^{d}}\hat{\nu}(z,k)\Big{[}\sum\limits_{j=1}^{p}e^{\frac{i\eta s}{\varepsilon}g(\xi_{j},k)}\psi(\xi_{j}-k)-\sum\limits_{l=1}^{q}e^{-\frac{i\eta s}{\varepsilon}g(\zeta_{l},-k)}\psi(\zeta_{l}+k)\Big{]}\frac{\mathrm{d}k}{(2\pi)^{d}}\,,$ (28)
where we explicitly indicate only the variables in $\psi$ that are shifted by ±$\pm k$.
@@
$\psi^{\theta}(z,v)=\sum\limits_{n\geq 0}\Big{(}\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\Big{)}^{n}\psi^{\theta}_{n}(z,v)\,,$ (29)
@@
where $[0,z]^{n}_{<}$ denotes the simplex $0\leq s_{1}\leq\cdots\leq s_{n}\leq z$. Moreover,
@@
Here, $A_{j}^{\prime}$ is a $(p+q)d\times Nd$ block diagonal matrix with uniformly bounded entries and $G_{j}^{\prime}$ is real valued $\forall j$.
@@
$\psi^{\theta}_{n}(z,v)=\int\limits_{[0,z]^{n}_{<}}\int\limits_{\mathbb{R}^{nd}}\Big{(}\prod\limits_{j=1}^{n}\hat{\nu}^{\theta}_{s_{j}}(k_{j})\Big{)}\sum\limits_{j=1}^{(p+q)^{n}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{n}\frac{\mathrm{d}k_{j}\mathrm{d}s_{j}}{(2\pi)^{d}}\,,$ (31)
@@
$A_{j}$ are $(p+q)d\times nd$ block matrices with $d\times d$ blocks such that $A_{j}$ has exactly one non zero entry per column block. This non zero entry equals 1 when it appears in the first $pd$ rows and equals $-1$ otherwise. $G_{j}$ can be defined recursively as follows:
Let $\epsilon_{m}=1$ if $m=1,\cdots,p$ and equal to $-1$ if $m=p+1,\cdots,p+q$. Also, let
$G_{1,m}(\vec{s},\vec{k},v)=s_{1}\epsilon_{m}g(v_{m},\epsilon_{m}k_{1})\,.$
@@
$G_{r+1,(p+q)(m^{\prime}-1)+m}(\vec{s},\vec{k},v)$ $=s_{r+1}\epsilon_{m}g(v_{m},\epsilon_{m}k_{r+1})$
@@
×$\times\sum\limits_{j=1}^{(p+q)^{2N}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{2N}\frac{\mathrm{d}k_{j}\mathrm{d}s_{j}}{(2\pi)^{d}}\,.$
Here, $P^{2N}$ denotes a permutation of pairings of $2N$ indices. The Dirac measure reduces integrals over $\vec{k}\in\mathbb{R}^{2Nd}$ to an integral over $\mathbb{R}^{Nd}$. This gives (30).
In (30), $A_{j}^{\prime}$ are $(p+q)d\times Nd$ block diagonal matrices of $d\times d$ blocks with either no non zero entries or exactly two non zero entries (either 1 or -1) per column block. Also, for a fixed arrangement $P^{2N}$ with $(l,m)\in P^{2N}$, $G_{j}^{\prime}(\vec{s},\vec{k},v)=G_{j}(\vec{s},B_{P^{2N}}(\vec{k}),v)$ where $B_{P^{2N}}(\vec{k})$ is a block matrix of size $2Nd\times 1$ with $k_{1},\cdots,k_{N}$ appearing as $l$th entries and $-k_{1},-k_{2},\cdots,-k_{N}$ appearing as corresponding $m$th entries. For example, when $N=2$, there are three possibilities for $P^{4}$: if $P^{4}=\{(1,2),(3,4)\}$, $B_{P^{4}}(\vec{k})=(k_{1},-k_{1},k_{2},-k_{2})^{\top}$ and if $P^{4}=\{(1,3),(2,4)\}$, $B_{P^{4}}(\vec{k})=(k_{1},k_{2},-k_{1},-k_{2})^{\top}$ and finally if $P^{4}=\{(1,4),(2,3)\}$, $B_{P^{4}}(\vec{k})=(k_{1},k_{2},-k_{2},-k_{1})^{\top}$. ∎

### v2

@@ Abstract
This paper concerns the propagation of high frequency wave-beams in highly turbulent atmospheres. Using a paraxial model of wave propagation, we show in the long-distance weak-coupling regime that the wavefields are approximately described by a complex Gaussian field whose scintillation index is unity. This provides a model of the speckle formation observed in many practical settings. The main step of the derivation consists in showing that closed-form moment equations in the Itô-Schrödinger regime are still approximately satisfied in the paraxial regime. This is done via Duhamel expansions, under the assumption that the random medium has Gaussian statistics. The rest of the proof is then an extension of results derived in [5].
@@ 1 Introduction
$\Big{(}\partial^{2}_{z}+\Delta_{x}+k_{0}^{2}n^{2}(z,x)\Big{)}p(z,x)={\color[rgb]{0,0,0}+}u_{0}(x)\delta_{0}{\color[rgb]{0,0,0}{}^{\prime}}(z)\,$, (1)
@@
with $0<\theta\ll 1$. As in [5], we consider the weak coupling regime $0<\varepsilon\ll 1$ in two sub-regimes, the kinetic regime when $\eta=1$ and the diffusive regime when $\eta\ll 1$. That the latter regime corresponds to long-distance propagation (optically thick media) may be obtained from the fact that distance of propagation times random strength (proportional to $\nu^{2}$) is $\eta/\varepsilon\times\varepsilon/\eta^{3}=\eta^{-2}\gg 1$. We will see below that $\varepsilon$ models the width of the incident beam $u_{0}$. We also provide a discussion of these scalings in application at the end of Section 3. For technical reasons as in [5], we need a large scale separation between $\varepsilon$ and $\eta$ in the diffusive regime and assume that $\eta=(\log|\log\varepsilon|)^{-1}$.
@@
The above equation is deduced from (4) by formally replacing $\theta^{-\frac{1}{2}}\nu(\frac{z}{\theta},x)\mathrm{d}z$ by its central limit approximation $\mathrm{d}B(z,x)$, where $B(z,x)$ is a mean-zero Gaussian process satisfying $\mathbb{E}[B(z,x)B(z^{\prime},y)]=\min(z,z^{\prime})R(x-y)$ with spatial covariance $R(x)$ defined in (10) below. The product $u^{\varepsilon}\mathrm{d}B$ should be interpreted in a standard Itô form [11, 12]. That the complex circular Gaussian conjecture holds for $u^{\varepsilon}$ in the limit $\varepsilon\to 0$ was shown recently in [5]. Careful moment estimates of $u^{\theta}$ allow us to extend such a result to the paraxial model of wave propagation.
@@ 2 Description of the wavefield, random medium, and moments
For simplicity, we set $2k_{0}=1$ since this (rescaled) parameter does not have any qualitative influence, and assume that the amplitude $\nu$ is appropriately rescaled so that the paraxial equation becomes
@@
$\partial_{z}\hat{u}^{\theta}=-\frac{i\eta}{\varepsilon}|\xi|^{2}\hat{u}^{\theta}+\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\int\limits_{\mathbb{R}^{d}}\hat{u}^{\theta}(z,\xi-k)\frac{{\color[rgb]{0,0,0}\hat{\nu}(\frac{\eta z}{\varepsilon\theta},\mathrm{d}k)}}{(2\pi)^{d}},\quad\hat{u}^{\theta}(0,\xi)=\hat{u}^{\theta}_{0}(\xi)\,$,
where we denote by $\hat{\nu}(z,\mathrm{d}k)$ the random measure such that $\nu(z,x)=\int_{\mathbb{R}^{d}}e^{ik\cdot x}\frac{\hat{\nu}(z,\mathrm{d}k)}{(2\pi)^{d}}$. For a collection of points $(X,Y)=(x_{1},\cdots,x_{p},y_{1},\cdots,y_{q})\in\mathbb{R}^{(p+q)d}$, define the $p+q$th moment of the wavefield $u^{\theta}$ at fixed $z$ by
@@
Gaussian random potential
@@
We assume that $\hat{\mathfrak{C}}(z,\cdot)$ is integrable in $z$, smooth in $k$ and rapidly decaying for large values of $z$. For concreteness and to simplify some derivations, we assume that for $\langle k\rangle:=\sqrt{1+|k|^{2}}$, we have
$\color[rgb]{0,0,0}\int_{\mathbb{R}^{d+1}}\langle{k}\rangle^{{d_{0}+3}}\langle{s}\rangle|\hat{\mathfrak{C}}(s,k)|\mathrm{d}s\mathrm{d}k=\lambda_{1}(d_{0})<\infty$}\,, (9)
where we assume that $d_{0}\geq d$. Except for improved regularity results obtained when $d_{0}\geq d$ in Lemma 4.7 (and Lemma 4.2) and for better stochastic continuity properties in Theorem 3.3, there is no loss in generality to assume that $d_{0}=d$, the lateral spatial dimension. We assume the standard symmetries for correlation functions: $\mathfrak{C}(z,x)=\mathfrak{C}(-z,-x)$ and denote by $R(x)$ the lateral spatial covariance
@@
Note that the Fourier transform $\hat{R}(k)$ is the power spectral density of a stationary process and is non-negative from Bochner’s theorem. We next assume that $R\in{\mathsf{L}}^{1}(\mathbb{R}^{d})\cap{\mathsf{L}}^{\infty}(\mathbb{R}^{d})$ with the Fourier transform $\hat{R}(k)=\hat{R}(-k)\leq\hat{{\mathsf{R}}}(k)=\hat{{\mathsf{R}}}(|k|)$ for some radially symmetric function $\hat{{\mathsf{R}}}$.
@@
with the Fourier transform $\hat{\mathfrak{C}}^{\theta}(s,k)$ defined in a similar manner.
Assumptions on the source
@@
uniformly in $\varepsilon\in(0,1]$. Here and below denote by $\mathcal{M}_{B}(\mathbb{R}^{m})$, the Banach space of finite signed measures on $\mathbb{R}^{m}$, equipped with the total variation (TV) norm $\|\cdot\|$. For measures with a density with respect to the Lebesgue measure (such as $\hat{u}^{\theta}_{0}(k)dk$ above with density $\hat{u}^{\theta}_{0}(k)$) then that norm is the ${\mathsf{L}}^{1}(\mathbb{R}^{d})$ norm of that density. We use the notation $\|\cdot\|$ both for the TV norm of bounded Radon measures or for the operator norm of operators acting on such measures. The standard ${\mathsf{L}}^{p}$ norms are denoted by $\|\cdot\|_{p}$.
Spatially rescaled random vector
@@ 3 Main results
The limiting approximations for $\phi^{\theta}$ as $\theta\to 0$ in the kinetic and diffusive regimes are the same as for the Itô-Schrödinger model (as $\varepsilon\to 0$) treated in [5]. The limits are described by second-order moments, which we now recall.
@@
[…]
@@
[…]
@@
In the kinetic regime $\eta=1$, $\Phi^{\theta}(z,r,X)\Rightarrow\Phi(z,r,X)$ in distribution as $\theta\to 0$ where $\Phi(z,r,X)$ is a random vector with elements $\{\phi_{j}\}_{j=1}^{N}$ satisfying $\mathbb{E}\phi_{j}=M_{1,0}(z,r,x_{j})$ and ${\color[rgb]{0,0,0}\mathcal{Z}}=\Phi-\mathbb{E}\Phi$ a mean zero circular complex Gaussian random vector with elements $\{{\color[rgb]{0,0,0}\mathcal{Z}}_{j}\}_{j=1}^{N}$ satisfying
$\mathbb{E}[{\color[rgb]{0,0,0}\mathcal{Z}}_{j}{\color[rgb]{0,0,0}\mathcal{Z}}^{\ast}_{l}]=M_{1,1}(z,r,x_{j},x_{l})-M_{1,0}(z,r,x_{j})M_{0,1}(z,r,x_{l}),\quad\mathbb{E}[{\color[rgb]{0,0,0}\mathcal{Z}}_{j}{\color[rgb]{0,0,0}\mathcal{Z}}_{l}]=0\,.$ (20)
@@
$\sup\limits_{0\leq s\leq z}\mathbb{E}|\phi^{\theta}(s,r,x+h)-\phi^{\theta}(s,r,x)|^{2n}\leq C(n,\alpha,z){\color[rgb]{0,0,0}\|h\|_{2}}^{2n\alpha}$ (22)
for $\|h\|_{2}<1$ uniformly in $\theta\in(0,\frac{1}{2}]$. Here, $\alpha$ is any scalar such that $\alpha\in(0,1]$ in the kinetic regime and $\alpha\in(0,1)$ in the diffusive regime. Let $n$ be such that $2n\alpha\geq d+2n\alpha_{-}$, $\alpha_{-}\in(0,\alpha)$. Then the process $\phi^{\theta}(z,r)$ is tight on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ and there exists a Hölder continuous version of $\phi^{\theta}(z,r)$ on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$.
@@
For $z$ and $r$ fixed and under the hypotheses of Theorem 3.3, the process $\phi^{\theta}(z,r,\cdot)$ converges in distribution as probability measures on $C^{0,\alpha_{-}}(\mathbb{R}^{d})$ to its unique (Gaussian) limit. This is characterized by limits of finite dimensional distributions shown to converge in Theorems 3.1 and 3.2 in the kinetic and diffusive regimes.
@@
We also note the independence of the macroscopic wavefields at different $r$ under the diffusive scaling in the following theorem.
Theorem 3.6.
For fixed z and $r\neq r^{\prime}$ in the diffusive regime, the random vectors $\Phi^{\theta}(z,r,X)$ and $\Phi^{\theta}(z,r^{\prime},X^{\prime})$ become statistically independent as $\theta\to 0$.
Scaling in applications
We briefly revisit the scaling (2) and its practical relevance, say for laser beam propagation through turbulence. There, $x$ was measured in terms of the smallest scale of turbulence, let us call it $l_{0}$, typically taken in practice to be in the $mm$ range. This means that medium fluctuations should be understood as $\nu=\nu\big{(}\frac{z}{l_{0}},\frac{x}{l_{0}}\big{)}$. With $Z$ a dimension-less parameter characterizing longitudinal distance of propagation, $w_{0}$ a parameter describing the width of the source, and $\sigma$ a covariance parameter for the random fluctuations $\nu=\sigma\tilde{\nu}$, we have the scaling and paraxial model after neglecting backscattering:
@@
A typical wavelength used in practice is $\lambda\approx 0.6\times 10^{-6}m$ and corresponds to $k_{0}\approx 10^{7}m^{-1}$. Under weak turbulence assumptions [10, 24, 1], the refractive index is approximated as $n(\vec{r})\approx 1+79\times 10^{-6}\frac{P(\vec{r})}{T(\vec{r})}$, where $P$ is the pressure in millibars and $T$ the temperature in Kelvin. Assuming the variations in the refractive index are mainly from temperature fluctuations ($T(\vec{r})=T+\delta T(\vec{r})$), this corresponds to $\delta n\approx 79\times 10^{-6}\frac{P}{T^{2}}\delta T$. So for changes in the temperature in the $0.1^{\circ}-1^{\circ}$ range, $\sigma=(\mathbb{E}[\delta n^{2}])^{\frac{1}{2}}\approx 10^{-7}-10^{-6}$. We also set the fluctuation lengthscale $l_{0}=2\times 10^{-3}m$ so that $\theta=5\times 10^{-5}$.
We thus observe that we reach the kinetic regime $\eta=1$ for a propagation distance of order $l_{0}Z=500m$ when $\sigma=10^{-7}$ with $\varepsilon=\frac{\eta k_{0}l_{0}}{Z}=0.08\ll 1$ when $w_{0}\geq\frac{10^{-3}}{\varepsilon}$, which is the case for the beam widths of a (quite a) few millimeters to meters found in applications [27, 30, 29, 1].
The diffusive regime is reached for $\eta\ll 1$, corresponding to $\sigma\approx 5\times 10^{-7}$ and $l_{0}Z$ in the range $10^{3}m$ typical in applications. For $l_{0}Z=10^{3}m$ so that $\eta=0.1$ and $\varepsilon=10^{-2}$ for concreteness, the proposed theory has a domain of validity for beam width $w_{0}=10^{-3}/\varepsilon$, i.e., for beams with a diameter in tens of centimeters to meter scale.
@@ 4.1 Phase compensated wave field and preparatory calculations
$\partial_{z}\tilde{u}^{\theta}=\frac{i}{(\eta\varepsilon\theta)^{\frac{1}{2}}}\int\limits_{\mathbb{R}^{d}}\tilde{u}^{\theta}(z,\xi-k)e^{\frac{i\eta z}{\varepsilon}g(\xi,k)}\frac{{\color[rgb]{0,0,0}\hat{\nu}(\frac{\eta z}{\varepsilon\theta},\mathrm{d}k)}}{(2\pi)^{d}},\quad\tilde{u}^{\theta}(0,\xi)=\hat{u}^{\theta}_{0}(\xi)\,$, (25)
@@
$L^{\theta}(s,\nu_{z})\psi=\int\limits_{\mathbb{R}^{d}}\Big{[}\sum\limits_{j=1}^{p}e^{\frac{i\eta s}{\varepsilon}g(\xi_{j},k)}\psi(\xi_{j}-k)-\sum\limits_{l=1}^{q}e^{-\frac{i\eta s}{\varepsilon}g(\zeta_{l},-k)}\psi(\zeta_{l}+k)\Big{]}\frac{{\color[rgb]{0,0,0}\hat{\nu}(z,\mathrm{d}k)}}{(2\pi)^{d}}\,$, (28)
where we explicitly indicate only the variables in $\psi$ that are shifted by $\pm k$.
@@
$\psi^{\theta}(z,v)=\sum\limits_{n\geq 0}\frac{{\color[rgb]{0,0,0}1}}{(\eta\varepsilon\theta)^{\frac{n}{2}}}\psi^{\theta}_{n}(z,v)\,$, (29)
@@
where $0,z]^{n}_{<}$ denotes the simplex $0\leq s_{1}\leq\cdots\leq s_{n}\leq z$. Moreover,
@@
Here, $A_{j}^{\prime}$ is a $(p+q)d\times Nd$ block diagonal matrix with uniformly bounded entries and $G_{j}^{\prime}$ is real valued $\forall j$. $P^{2N}$ denotes a permutation of pairings of $2N$ indices.
@@
$\psi^{\theta}_{n}(z,v)=\int\limits_{[0,z]^{n}_{<}}\int\limits_{\mathbb{R}^{nd}}\sum\limits_{j=1}^{(p+q)^{n}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{n}{\color[rgb]{0,0,0}\frac{\hat{\nu}^{\theta}_{s_{j}}(\mathrm{d}k_{j})}{(2\pi)^{d}}\mathrm{d}s_{j}}\,$, (31)
@@
$A_{j}$ are $(p+q)d\times nd$ block matrices with $d\times d$ blocks such that $A_{j}$ has exactly one non zero entry per column block. This non zero entry equals 1 when it appears in the first $pd$ rows and equals 1-1 otherwise. $G_{j}$ can be defined recursively as follows:
Let $\epsilon_{m}=1$ if $m=1,\cdots,p$ and equal to 1-1 if $m=p+1,\cdots,p+q$. Also, let
$G_{1,m}(\vec{s},\vec{k},v)=s_{1}\epsilon_{m}g(v_{m},\epsilon_{m}k_{1}){\color[rgb]{0,0,0}+\epsilon_{m}\frac{\pi}{2}\frac{\varepsilon}{\eta}}\,.$
@@
$G_{r+1,(p+q)(m^{\prime}-1)+m}(\vec{s},\vec{k},v)$ $=s_{r+1}\epsilon_{m}g(v_{m},\epsilon_{m}k_{r+1}){\color[rgb]{0,0,0}+\epsilon_{m}\frac{\pi}{2}\frac{\varepsilon}{\eta}}$
@@
$\times\sum\limits_{j=1}^{(p+q)^{2N}}e^{\frac{i\eta}{\varepsilon}G_{j}(\vec{s},\vec{k},v)}\psi^{\theta}(0,v-A_{j}\vec{k})\prod\limits_{j=1}^{2N}\frac{\mathrm{d}k_{j}\mathrm{d}s_{j}}{(2\pi)^{d}}\,.$
The Dirac measure reduces integrals over $\vec{k}\in\mathbb{R}^{2Nd}$ to an integral over $\mathbb{R}^{Nd}$. This gives (30).
In (30), $A_{j}^{\prime}$ are $(p+q)d\times Nd$ block diagonal matrices of $d\times d$ blocks with either no non zero entries or exactly two non zero entries (either 1 or -1) per column block. Also, for a fixed arrangement $P^{2N}$ with $(l,m)\in P^{2N}$, $G_{j}^{\prime}(\vec{s},\vec{k},v)=G_{j}(\vec{s},B_{P^{2N}}(\vec{k}),v)$ where $B_{P^{2N}}(\vec{k})$ is a block matrix of size $2Nd\times 1$ with $k_{1},\cdots,k_{N}$ appearing as $l$th entries and $k_{1},-k_{2},\cdots,-k_{N}$ appearing as corresponding $m$th entries. For example, when $N=2$, there are three possibilities for $P^{4}$: if $P^{4}=\{(1,2),(3,4)\}$, $B_{P^{4}}(\vec{k})=(k_{1},-k_{1},k_{2},-k_{2})^{\top}$ and if $P^{4}=\{(1,3),(2,4)\}$, $B_{P^{4}}(\vec{k})=(k_{1},k_{2},-k_{1},-k_{2})^{\top}$ and finally if $P^{4}=\{(1,4),(2,3)\}$, $B_{P^{4}}(\vec{k})=(k_{1},k_{2},-k_{2},-k_{1})^{\top}$. ∎